use the discriminant to determine the number of real roots of each equation and then solve each equation using the quadratic formula.
2t^2 + 1 = 6t

Answer by TimothyLamb(4379)

2t^2 + 1 = 6t
2t^2 - 6t + 1 = 0
---
the above quadratic equation is in standard form, with a=2, b=-6, and c=1
---
from the quadratic formula the discriminant is: b^2 - 4ac
---
b^2 - 4ac
(-6)^2 - 4(2)(1)
36 - 8 = 28
---
Answer 1:
since the discriminant > 0 the function has two real roots
---
to solve the above equation using the quadratic formula, plug this:
2 -6 1
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 2:
the roots (solutions) of the above quadratic equation are:
2.82287566
0.177124344
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php